Question 15 Type numbers in the boxes. Find the p-value for the hypothesis test. A random sample of size 51 is taken. The samround to 4 decimals
Answer #1

Solution :

\mu= 400

\bar x =392

\sigma =83

n = 51

This is the left tailed test .

The null and alternative hypothesis is ,

H0 :  \mu = 400

Ha : \mu <400

Test statistic = z

= (\bar x - \mu ) / \sigma / \sqrt n

= (392-400) / 83 / \sqrt 51

= −0.69

P(z <-0.69 ) = 0.2451

P-value = 0.2451